# Inferring collective synchrony observing spiking of one or several neurons

### v1

@@ I Introduction
Quantifying collective synchrony is a common task in the physics of complex systems in general, not only in neuroscience. In many situations, the solution is well-known and straightforward. So, if the phases of all interacting units are known, the Kuramoto order parameter - a quantity between zero and one - gives the desired answer. An alternative way is to compute the standard deviation of the collective mode; in contradistinction to the Kuramoto order parameter, this quantity is not normalized. However, it allows one to trace the synchronization transition if one monitors the collective mode’s standard deviation while varying the degree of interaction between the units. The problem becomes much less trivial if we observe only a small fraction of the population, possibly even one unit. In our recent publication Pikovsky and Rosenblum (2024), we suggested a solution to this problem, demonstrating quantification of synchrony from a partial observation under certain conditions. However, the approach of Ref. Pikovsky and Rosenblum (2024) works with oscillators generating smooth signals, while in neuroscience applications, one often deals with spike trains that can be considered as point processes. This paper tackles this problem and develops a technique for quantifying synchrony in a large highly interconnected neuronal network from the recordings of spiking events of one or several units. It is known that such networks can be treated in the mean-field approximation, i.e., for simplicity, one can assume the global (all-to-all) coupling.
@@ II.1.1 Noisy spiking and bursting Hindmarsh-Rose neurons
Figure 1: The top panel illustrates the spiking of a noisy Hindmarsh-Rose neuron, while the bottom panel depicts the bursting of HR neurons for $I_{n}=2.95$, $I_{n}=3$, and $I_{n}=3.05$. (The second and third curves are shifted upwards for better visibility.) From each trace, the point process is generated by taking the instants of the threshold crossing from below, the dotted line $x=1$ marks the threshold.
Figure 2: The top panel illustrates the synchronization transition in a system of $N=20000$ coupled identical noisy Hindmarh-Rose neurons, see Eq. (1); here, we plot the mean field variance as a function of the coupling strength $\varepsilon$. The mean field $X_{M}$ is computed from $M\leq N$ neurons. (Notice the logarithmic scale of the vertical axis.) The plot demonstrates that observation of only $100$ units out of $N$ provides a reliable indication of synchrony. The middle panels present the raster plots; here, the instants of spiking are marked by a dot (spikes from only 500 neurons are shown for better visibility). The raster plots are given for four values of the coupling strength $\varepsilon$ marked by dotted vertical lines in the top panel ($0.25$, $0.35$, $0.5$, $0.7$). The bottom panels exhibit the corresponding time traces of the mean field $X=X_{N}$.
@@
A single neuron ($\varepsilon=0$) is producing an irregular, due to noise in the setup 1 or due to chaos in the setup 2, sequence of spikes or bursts, as illustrated in Fig. 1. A population of neurons, with increase of the coupling strength $\varepsilon$, experiences a transition to collective synchrony, as illustrated in Fig. 2 for noisy identical spiking neurons and in Fig. 3 for nonidentical chaotic bursting neurons. The transition can be determined by following the variance of the mean field $X(t)$. In the thermodynamic limit $N\to\infty$, in the asynchronous regime this mean field is constant, and it is oscillatory beyond the synchronization transition. Using the variance of the mean field as an indicator for synchronization transition has been suggested already in early studies of coupled disordered (noisy or chaotic) systems Ginzburg and Sompolinsky (1994); Hansel and Sompolinsky (1996); Pikovsky, Rosenblum, and Kurths (1996); Golomb, Hansel, and Mato (2001). Moreover, one can use the partial mean field $X_{M}$ averaged not over the whole population, but over a subset of $M<N$ units Ginzburg and Sompolinsky (1994); Hansel and Sompolinsky (1996); Golomb, Hansel, and Mato (2001). One expects that in the asynchronous case the variance is $\sim M^{-1}$ due to finite-size effects, so for sufficiently large $M$ the transition should be visible. Indeed, as can be seen in Fig. 2, the variance of a partial field observed from $M=100$ neurons is almost as good an indicator of synchrony as the variance computed for $M=N=20000$. Obviously, the variance of $X_{1}$, i.e., of the trace of an individual unit, is practically constant and does not provide information about the transition. However, we demonstrate below that we can reveal synchrony by observing only one unit.
@@ II.1.2 Brunel-Hakim model for sparse synchrony
In a series of papers Brunel and Hakim (1999, 2008); Ostojic, Brunel, and Hakim (2009), Brunel and Hakim suggested and explored a simple model exhibiting nontrivial dynamics where the mean-field frequency is much higher than the firing frequency of individual units; they picked the term sparse synchrony to emphasize this property. The model comprises noisy leaky integrate-and-fire neurons (LIF) $v_{n}(t)$ coupled via global fields $X,Y$:
@@
Here $\xi_{n}(t)$ are independent Gaussian white noises with zero mean and unit intensity. As it is typical for the integrate-and-fire models, a hybrid system (2) describes the evolution of the voltage $v_{n}$ of the neuron $n$ until the voltage achieves a threshold $v_{u}$; at this moment denoted as $t_{k}^{(n)}$, the neuron $n$ fires an action potential and is instantaneously reset to the level $v_{d}$. The produced spike contributes to the instantaneous change of the global field $Y(t)$. As in the case of HR neurons, we denote the coupling strength $\varepsilon$. Due to noise, uncoupled noisy LIF neurons produce irregular spike trains. With the increase of $\varepsilon$, one observes a transition from the disordered state, where fields $X,Y$ are constants (in the thermodynamic limit), to collective synchrony, where the fields $X,Y$ oscillate, see Fig. 4222A detailed study shows that the system exhibits bistability and hysteresis in the transition; however, these properties are irrelevant for our study. To avoid ambiguity in the results, in the simulations we always use the nearly-asynchronous initial conditions. Remarkably, the period of these oscillations is much smaller than the characteristic interspike interval of a single neuron. Thus, this regime was termed “sparsely synchronized neuronal oscillations” in Brunel and Hakim (2008). In the simulation, we use the following parameter values: $v_{u}=5$, $v_{d}=14$, $\tau_{m}=5$, $\tau_{d}=6$, $\tau_{r}=1$, $\sigma=0.5$, and $I_{0}=50$.
@@ II.2.1 Observation of many neurons
First, we mention a relatively simple case where one observes all or many neurons from the ensemble. The standard approach presents all observations as a raster plot, where each spike is represented with a dot marker and all available spike trains are superimposed, cf. Fig. 2. Collective synchrony in such a representation appears as a pronounced modulation of the time-dependent density of markers, while in the absence of synchrony, this density is constant. Thus, the collective synchrony can be inferred by calculating the time-dependent density (instantaneous population firing rate) from the raster plots; a measure of the macroscopic density variation characterizes the synchrony level Singer (1999); Ostojic, Brunel, and Hakim (2009); Ciba et al. (2018).
@@ II.2.4 Quantification of a regular component in the covariance density
One possible choice for $Q^{u}$ would be $Q^{u}(p)=\tau_{(1)}+p(\tau_{(L)}-\tau_{(1)})$. However, such a choice is sensitive to boundary effects. Thus, we suggest to apply a linear fit to the set $\tau_{(m)}$, i.e., to approximate it as $\tau_{(m)}\approx a+bm$ (for a standard procedure for the linear fit computation, see, e.g., in Press et al. (1992)). Fitting corresponds to taking the uniform quantile function as $Q^{u}(p)=a+bLp$. Substitution of this function and the empirical cumulative covariance distribution function (9) (10) leads, after representing the integral as a sum, to the following simple formula for the KRW distance
@@
Since $D$ quantifies a regular component in the dynamics, it serves as an order parameter for the transition to synchrony.  [^3] One can also consider generalized KRW distances by changing the definition of the norm in (13); in general, one could calculate a $q$-norm as $(\sum_{m}|C(\tau_{(m)})|^{q})^{1/q}$. However, only for $q=1$ the simple representation (11) is valid. Of course, this parameter does not vanish in the asynchronous case for finite samples, but for large $L$ its value can be relatively small Del Barrio, Giné, and Matrán (1999). We emphasize the similarity between the order parameters $D$ (Eq. (13)) and $W$ (Eq. (4)): both calculations involve two constants. However, while in the case of continuous signals, constants $\Theta_{1,2}$ explicitly enter Eq. (4), for the point processes, constants $\theta_{1,2}$ appear implicitly in the selection of the time differences $\tau_{m}$ according to the condition (8).
@@ IV Discussion and conclusion
We demonstrated that an appropriate processing of one or several units’ activity reveals a macroscopic rhythm in a neuronal population. This result aligns with the findings that a small subset of neuronal populations is involved in the visual perception of complex images (Quian Quiroga, Reddy, and Kreiman, 2005) or conveys information sufficient for neuroprosthetics purposes (Willett et al., 2023). The developed technique quantifies the level of synchrony in the neuronal population from the observed spiking events obtained via threshold-crossing. The height of the spikes is not accounted for, so the input data are point processes. The analysis of the covariance density is mathematically equivalent to the analysis of the Bartlett spectral measure. Indeed, the transition to synchrony is reflected by the appearance of a discrete spectral component (in the limit $N\to\infty$) what corresponds to non-decaying $C(\tau)$. Thus, the measure $D$ quantifies the discrete spectral component, and this computation is analogous to Wiener’s lemma, which yields a similar quantity for continuous processes from the autocovariance function.
@@
The simple and computationally inexpensive algorithm provides a single number, which we denote as the KRW distance $D$. This order parameter successfully distinguishes synchronous and asynchronous states. The drawback of this measure is that it is not normalized; hence, a single computation does not say whether $D$ is large (synchrony) or small (asynchrony). In this case, one has to plot the empirical cumulative covariance distribution function $C(\tau)$ and inspect it visually. Another ad hoc solution is to rely on the surrogate test. However, if observations of synchronous and asynchronous states are given, the corresponding values of $D$ differ by more than one order of magnitude. A clear advantage of the approach is its stability regarding imperfect measurement – missing about 10 or 15 percent of spikes do not affect the result. Another beneficial feature is that the number of neurons contributing to the observed point process is irrelevant, and spike sorting is not needed. The suggested technique works equally well with the data from one neuron and a mixed point process containing spikes from many neurons. Actually, in the last case, the performance is even better (though the number of these neurons remains much smaller than the population size). These features make the developed approach an efficient tool for experimental and model studies.
The exploited test models assumed that synchrony emerges due to coupling within a population. In reality, neuronal populations are typically not isolated but affected by other populations and/or external fields. In all these cases, our approach will reveal and quantify synchronous firing, though it cannot help to establish the origin of synchrony.
Acknowledgements We thank A. Neiman and V. Nikulin for useful comments.

### v2

@@ I Introduction
Quantifying collective synchrony is a common task in the physics of complex systems in general, not only in neuroscience. In many situations, the solution is well-known and straightforward. So, if the phases of all interacting units are known, the Kuramoto order parameter - a quantity between zero (asynchrony) and one (complete synchrony) - gives the desired answer. An alternative approach is to compute the collective mode’s standard deviation; unlike the Kuramoto order parameter, this is not a dimensionless quantity. Thus, one cannot estimate the synchrony level when computing it for a certain network’s state. However, this quantity allows one to trace the synchronization transition if one monitors the collective mode’s standard deviation and observes its essential increase while varying the degree of interaction between the units. 111 For supplementary information, one can compare the standard deviation of the collective mode with that of the individual unit’s oscillation: in the case of synchrony, these quantities are of the same order. However, this comparison requires additional measurements.
The problem becomes much less trivial if we observe only a small fraction of the population, possibly even one unit. In our recent publication Pikovsky and Rosenblum (2024), we suggested a solution to this problem, demonstrating quantification of synchrony from a partial observation under certain conditions. However, the approach of Ref. Pikovsky and Rosenblum (2024) works with oscillators generating smooth signals, while in neuroscience applications, one often deals with spike trains that can be considered as point processes. This paper tackles this problem and develops a technique for quantifying synchrony in a large highly interconnected neuronal network from the recordings of spiking events of one or several units. It is known that such networks can be treated in the mean-field approximation, i.e., for simplicity, one can assume the global (all-to-all) coupling.
@@ II.1.1 Noisy spiking and bursting Hindmarsh-Rose neurons
Figure 1: The panel (a) illustrates the spiking of a noisy Hindmarsh-Rose neuron, while the panel (b) depicts the bursting of HR neurons for $I_{n}=2.95$, $I_{n}=3$, and $I_{n}=3.05$. (The second and third curves are shifted upwards for better visibility.) From each trace, the point process is generated by taking the instants of the threshold crossing from below, the dotted line $x=1$ marks the threshold. Notice the difference in time scales in (a) and (b).
@@
Figure 2: The top panel illustrates the synchronization transition in a system of $N=20000$ coupled identical noisy Hindmarsh-Rose neurons, see Eq. (1); here, we plot the mean field variance as a function of the coupling strength $\varepsilon$. The mean field $X_{M}$ is computed from $M\leq N$ neurons. (Notice the logarithmic scale of the vertical axis.) The plot demonstrates that observation of only $100$ units out of $N$ provides a reliable indication of synchrony. The middle panels present the raster plots; here, the instants of spiking are marked by a dot (spikes from only 500 neurons are shown for better visibility). The raster plots are given for four values of the coupling strength $\varepsilon$ marked by dotted vertical lines in the top panel ($0.25$, $0.35$, $0.5$, $0.7$). The bottom panels exhibit the corresponding time traces of the mean field $X=X_{N}$.
@@
A single neuron ($\varepsilon=0$) is producing an irregular, due to noise in the setup 1 or due to chaos in the setup 2, sequence of spikes or bursts, as illustrated in Fig. 1. A population of neurons, with increase of the coupling strength $\varepsilon$, experiences a transition to collective synchrony, as illustrated in Fig. 2 for noisy identical spiking neurons and in Fig. 3 for nonidentical chaotic bursting neurons. The transition can be determined by following the variance of the mean field $X(t)$. In the thermodynamic limit $N\to\infty$, in the asynchronous regime this mean field is constant, and it is oscillatory beyond the synchronization transition. Using the variance of the mean field as an indicator for synchronization transition has been suggested already in early studies of coupled disordered (noisy or chaotic) systems Ginzburg and Sompolinsky (1994); Hansel and Sompolinsky (1996); Pikovsky et al (1996); Golomb et al (2001). Moreover, one can use the partial mean field $X_{M}$ averaged not over the whole population, but over a subset of $M<N$ units Ginzburg and Sompolinsky (1994); Hansel and Sompolinsky (1996); Golomb et al (2001). One expects that in the asynchronous case the variance is $\sim M^{-1}$ due to finite-size effects, so for sufficiently large $M$ the transition should be visible. Indeed, as can be seen in Fig. 2, the variance of a partial field observed from $M=100$ neurons is almost as good an indicator of synchrony as the variance computed for $M=N=20000$. Obviously, the variance of $X_{1}$, i.e., of the trace of an individual unit, is practically constant and does not provide information about the transition. However, we demonstrate below that we can reveal synchrony by observing only one unit.
@@ II.1.2 Brunel-Hakim model for sparse synchrony
In a series of papers Brunel and Hakim (1999, 2008); Ostojic et al (2009), Brunel and Hakim suggested and explored a simple model exhibiting nontrivial dynamics where the mean-field frequency is much higher than the firing frequency of individual units; they picked the term sparse synchrony to emphasize this property. The model comprises noisy leaky integrate-and-fire neurons (LIF) $v_{n}(t)$ coupled via global fields $X,Y$:
@@
Here $\xi_{n}(t)$ are independent Gaussian white noises with zero mean and unit intensity; $\tau_{m}$ is the membrane time constant, and $\tau_{r}$ and $\tau_{d}$ are the decay and rise time of the postsynaptic current. As it is typical for the integrate-and-fire models, a hybrid system (2) describes the evolution of the voltage $v_{n}$ of the neuron $n$ until the voltage achieves a threshold $v_{u}$; at this moment denoted as $t_{k}^{(n)}$, the neuron $n$ fires an action potential and is instantaneously reset to the level $v_{d}$. The produced spike contributes to the instantaneous change of the global field $Y(t)$. As in the case of HR neurons, we denote the coupling strength $\varepsilon$. The equations for global fields $X,Y$ represent synaptic filtering.
Due to noise, uncoupled noisy LIF neurons produce irregular spike trains. With the increase of $\varepsilon$, one observes a transition from the disordered state, where fields $X,Y$ are constants (in the thermodynamic limit), to collective synchrony, where the fields $X,Y$ oscillate, see Fig. 4333A detailed study shows that the system exhibits bistability and hysteresis in the transition; however, these properties are irrelevant for our study. To avoid ambiguity in the results, in the simulations we always use the nearly-asynchronous initial conditions. Remarkably, the period of these oscillations is much smaller than the characteristic interspike interval of a single neuron. Thus, this regime was termed “sparsely synchronized neuronal oscillations” in Brunel and Hakim (2008). In the simulation, we use the following parameter values: $v_{u}=5$, $v_{d}=14$, $\tau_{m}=5$, $\tau_{d}=6$, $\tau_{r}=1$, $\sigma=0.5$, and $I_{0}=50$.
@@ II.2.1 Observation of many neurons
First, we mention a relatively simple case where one observes all or many neurons from the ensemble. The standard approach presents all observations as a raster plot, where each spike is represented with a dot marker and all available spike trains are superimposed, cf. Fig. 2. Collective synchrony in such a representation appears as a pronounced modulation of the time-dependent density of markers, while in the absence of synchrony, this density is constant. Thus, the collective synchrony can be inferred by calculating the time-dependent density (instantaneous population firing rate) from the raster plots; a measure of the macroscopic density variation characterizes the synchrony level Singer (1999); Ostojic et al (2009); Ciba et al (2018).
@@ II.2.4 Quantification of a regular component in the covariance density
One possible choice for $Q^{u}$ would be $Q^{u}(p)=\tau_{(1)}+p(\tau_{(L)}-\tau_{(1)})$. However, such a choice is sensitive to boundary effects. Thus, we suggest to apply a linear fit to the set $\tau_{(m)}$, i.e., to approximate it as $\tau_{(m)}\approx a+bm$ (for a standard procedure for the linear fit computation, see, e.g., in Press et al (1992)). Fitting corresponds to taking the uniform quantile function as $Q^{u}(p)=a+bLp$. Substitution of this function and the empirical cumulative covariance distribution function (9) (10) leads, after representing the integral as a sum, to the following simple formula for the KRW distance
@@
Since $D$ quantifies a regular component in the dynamics, it serves as an order parameter for the transition to synchrony.  [^4] One can also consider generalized KRW distances by changing the definition of the norm in (13); in general, one could calculate a $q$-norm as $(\sum_{m}|C(\tau_{(m)})|^{q})^{1/q}$. However, only for $q=1$ the simple representation (11) is valid. Of course, this parameter does not vanish in the asynchronous case for finite samples, but for large $L$ its value can be relatively small Del Barrio et al (1999). We emphasize the similarity between the order parameters $D$ (Eq. (13)) and $W$ (Eq. (4)): both calculations involve two constants. However, while in the case of continuous signals, constants $\Theta_{1,2}$ explicitly enter Eq. (4), for the point processes, constants $\theta_{1,2}$ appear implicitly in the selection of the time differences $\tau_{m}$ according to the condition (8).
@@ IV Discussion and conclusion
We demonstrated that an appropriate processing of one or several units’ activity reveals a macroscopic rhythm in a neuronal population. This result aligns with the findings that a small subset of neuronal populations is involved in the visual perception of complex images (Quian Quiroga et al, 2005) or conveys information sufficient for neuroprosthetics purposes (Willett et al, 2023). The developed technique quantifies the level of synchrony in the neuronal population from the observed spiking events obtained via threshold-crossing. The height of the spikes is not accounted for, so the input data are point processes. The analysis of the covariance density is mathematically equivalent to the analysis of the Bartlett spectral measure. Indeed, the transition to synchrony is reflected by the appearance of a discrete spectral component (in the limit $N\to\infty$) what corresponds to non-decaying $C(\tau)$. Thus, the measure $D$ quantifies the discrete spectral component, and this computation is analogous to Wiener’s lemma, which yields a similar quantity for continuous processes from the autocovariance function.
@@
The simple and computationally inexpensive algorithm provides a single number, which we denote as the KRW distance $D$. This order parameter successfully distinguishes synchronous and asynchronous states. The drawback of this measure is that it is not normalized; hence, a single computation does not say whether $D$ is large (synchrony) or small (asynchrony). In this case, one has to plot the empirical cumulative covariance distribution function $C(\tau)$ and inspect it visually. Another ad hoc solution is to rely on the surrogate test. However, if observations of synchronous and asynchronous states are given, the corresponding values of $D$ differ by more than one order of magnitude. A clear advantage of the approach is its stability regarding imperfect measurement – missing about 10 or 15 percent of spikes do not affect the result. Another beneficial feature is that the number of neurons contributing to the observed point process is irrelevant, and spike sorting is not needed. The suggested technique works equally well with the data from one neuron and a mixed point process containing spikes from many neurons. Actually, in the last case, the performance is even better (though the number of these neurons remains much smaller than the population size). These features make the developed approach an efficient tool for experimental and model studies. This paper presented examples where we considered statistically stationary situations only. Stationarity allowed for analyzing a rather long time series, for which regularity detection in spike trains works reliably. In a non-stationary case, one has to perform the analysis within a finite time window, i.e. with a rather small time series. As we have demonstrated in Fig. 14, for a short sequence of spikes the ECCDF $C(\tau)$ becomes rather noisy. The data of Fig. 14 suggest that at least several hundred spikes are necessary for regularity detection. A more detailed statistical analysis is a subject of future studies.
This paper focuses on characterizing regularity in the observed point process or several point processes. In some cases, continuous-time measurements of the Local Field Potentials (LFP) can be performed simultaneously with the registration of the spiking activity. If LFP observations are available, one can use them to characterize the regularity of the neuronal population using well-established methods for continuous processes. The correspondence between the information provided by the LFP analysis and our technique needs further clarification. It is known that the variation of the LFP reflects spiking multi-unit activity, but their interrelationship is complicated and depends on the frequency range (Telenczuk and Destexhe, 2022). Since the LFP-spike interrelation is very sensitive to neuronal correlations (Telenczuk and Destexhe, 2022), we expect that for strong synchrony, the LFP analysis and point process processing provide similar conclusions, while for low and intermediate synchrony levels, our technique may be more informative. Additionally, the LFP is a broad-band signal, and the results of its spectral-based quantification depend on the choice of the frequencies of interest and preprocessing techniques, such as filtration. We also mention that point process analysis admits a simple surrogate test for significance (reshuffling of inter-spike intervals), while the corresponding tests for continuous-time processes are problematic.
Finally, while the regularity of spiking is usually associated with synchrony, the sources of the synchrony can potentially be different. In the examples explored in this paper, synchrony appears due to the interaction of the neurons within the population. However, a similar state can appear due to regular external action on the neurons from other brain regions or external fields. Clearly, the method of regularity detection proposed in this paper cannot resolve the source of regularity; however, it can be a part of a more comprehensive analysis of larger brain areas.
We thank A. Neiman and V. Nikulin for useful comments.